OpenStudy (anonymous):

a car is traveling in a straight line and decelerating at a constant rate. after 3.2 seconds it has traveled 88m and slowed to a speed of 27 m/s what is the initial speed i got 52 m/s

OpenStudy (anonymous):

\[v^2-2as=u^2\]

OpenStudy (perl):

what does s and u stand for?

OpenStudy (anonymous):

Put \[a=\frac{v-u}{t}\]

OpenStudy (anonymous):

s distance u initial velocity

OpenStudy (anonymous):

and the plug a back into the equation that is what i did but im not sure if my math is right

OpenStudy (anonymous):

other people had like 97m/s which didnt make sense to me

OpenStudy (perl):

v is final velocity, u is initial velocity

OpenStudy (anonymous):

Better use \[s=ut+\frac{1}{2}at^2\]

OpenStudy (perl):

so youre using kinematic equations under the assumption that acceleration is constant

OpenStudy (perl):

might as well write them all out :)

OpenStudy (anonymous):

Help me

OpenStudy (perl):

youre doing good

OpenStudy (perl):

there are 4 kinematic equations under constant acceleration

OpenStudy (anonymous):

Gotto go i told everything would need to do this problem

OpenStudy (anonymous):

i used x=xknot+vknot(t)+.5a(t)^2 where x= position, v=velocity a=acceleration and t= time

OpenStudy (anonymous):

i solved for and the plugged it into v=vknot+a(t) and solved for vknot

OpenStudy (anonymous):